# A transiting multi-planet system in the 61 million year old association Theia 116

### v1

@@ 6.1 Atmospheric Evolution
Planets with H/He dominated atmospheres are thought to be born larger, and through evolutionary processes, experience contraction and mass loss resulting in a decrease in the observed radius. At $\mathord{\sim}61$ Myr, TIC 434398831 b and c are likely still experiencing mass loss via atmospheric erosion (Owen & Wu, 2013; Owen & Schlichting, 2023; Gupta & Schlichting, 2019; Rogers & Owen, 2021; Rogers et al., 2023). We investigate the eventual fates of TIC 434398831 b and c by simulating their radius evolution following the semi-analytic model for thermal and photoevaporative evolution from Owen & Wu (2017). We implement an MCMC using emcee to explore possible initial conditions and radius evolution of TIC 434398831 b and c. We allow the core mass $M_{b,\mathrm{core}}$, bulk core density $\rho_{b,\mathrm{core}}$, initial envelope mass fraction $f_{env,b}$, and constant mass loss efficiency $\eta_{b}$ to be free parameters. The initial envelope mass fraction is defined at 10 Myr (i.e. disk dissipation time scale) as per literature standards. We fix the stellar mass and orbital period with our best-fit values derived in Section 4. We anchor the planet radius and the system age, rejecting chains that do not result in an evolutionary path that falls within the $1\sigma$ range of our measured planet radii at our system age and associated uncertainties.
TIC 434398831 b currently falls into the typical mini-Neptune regime ($2-4\,R_{\oplus}$), however, its eventual fate is unknown, as it may evolve across the radius valley to be a super-Earth ($1-2\,R_{\oplus}$). Our simulations predict $M_{b,\mathrm{core}}=6.6_{-1.5}^{+1.1}\,M_{\oplus}$, $f_{env,b}=0.41_{-0.21}^{+0.24}$, $\eta_{c}=0.087_{-0.036}^{+0.073}$, and $\rho_{b,\mathrm{core}}=4.5^{+1.3}_{-1.3}\,\mathrm{g\,cm^{-3}}$. While the majority of our simulated evolutionary paths predict TIC 434398831 b to remain in the mini-Neptune regime, $\mathord{\sim}1\%$ of our simulations result in TIC 434398831 b being stripped of its volatile envelope and becoming a super-Earth. We present a sample of 30 evolutionary tracks for TIC 434398831 b in Figure 13.
@@
Both planets are predicted to have significant initial mass envelope fractions, $f_{env,b}=0.41$ and $f_{env,c}=0.65$. These values are significantly larger than the predicted average initial mass envelope fraction for the Kepler distribution, $\mathord{\sim}0.10$. However, JWST observations of the $\mathord{\sim}23$ Myr, V1298 Tau b revealed a H/He envelope fraction of $\mathord{\sim}40\%$ (Barat et al., 2024) similar to the predicted mass envelope fractions of TIC 434398831 b and c. While these simulations only account for mass-loss via photoevaporation, it is likely additional mass-loss mechanisms are at play and work in hand with photoevaporation to sculpt the distribution of the mature population (e.g. Gupta & Schlichting, 2019; Owen & Schlichting, 2023; Rogers et al., 2024).

### v2

@@ 6.1 Atmospheric Evolution
Planets with H/He dominated atmospheres are thought to be born larger, and through evolutionary processes, experience contraction and mass loss resulting in a decrease in the observed radius. At $\mathord{\sim}61$ Myr, TIC 434398831 b and c are likely still experiencing mass loss via atmospheric erosion (Owen & Wu, 2013; Owen & Schlichting, 2023; Gupta & Schlichting, 2019; Rogers & Owen, 2021; Rogers et al., 2023). We investigate the eventual fates of TIC 434398831 b and c by simulating their radius evolution following the semi-analytic model for thermal and photoevaporative evolution from Owen & Wu (2017). We implement an MCMC using emcee to explore possible initial conditions and radius evolution of TIC 434398831 b and c. We allow the core mass $M_{b,\mathrm{core}}$, bulk core density $\rho_{b,\mathrm{core}}$, initial envelope mass fraction $f_{env,b}$, and constant mass loss efficiency $\eta_{b}$ to be free parameters. The initial envelope mass fraction is defined at 10 Myr (i.e. disk dissipation time scale) as per literature standards. We fixed the stellar mass with our best-fit value derived in Section 4 and adopted the empirical relation presented in Jackson et al. (2012) to model the evolution of the high-energy luminosity of the host star. We make use of the MESA stellar evolution models interpolated with minimint using our best fit stellar mass and metallicity, at each step in our planetary evolution model, we adopt the corresponding stellar radius and effective temperature. We fixed the orbital period at our best-fit value, which is used to estimate the equilibrium temperature of the planet. We anchored the planet radius and the system age, rejecting chains that do not result in an evolutionary path that falls within the $1\sigma$ range of our measured planet radii at our system age and associated uncertainties.
TIC 434398831 b currently falls into the typical mini-Neptune regime ($2-4\,R_{\oplus}$), however, its eventual fate is unknown, as it may evolve across the radius valley to be a super-Earth ($1-2\,R_{\oplus}$). Our simulations predict $M_{b,\mathrm{core}}=6.8_{-1.5}^{+1.9}\,M_{\oplus}$, $f_{env,b}=0.41_{-0.21}^{+0.24}$, $\eta_{c}=0.087_{-0.036}^{+0.073}$, and $\rho_{b,\mathrm{core}}=4.5^{+1.3}_{-1.3}\,\mathrm{g\,cm^{-3}}$. While the majority of our simulated evolutionary paths predict TIC 434398831 b to remain in the mini-Neptune regime, $\mathord{\sim}2.5\%$ of our simulations result in TIC 434398831 b being stripped of its volatile envelope and becoming a super-Earth. We present a sample of 30 evolutionary tracks for TIC 434398831 b in Figure 14.
@@
Both planets are predicted to have significant initial mass envelope fractions, $f_{env,b}=0.41$ and $f_{env,c}=0.65$. These values are significantly larger than the predicted average initial mass envelope fraction for the Kepler distribution, $\mathord{\sim}0.10$. However, HST observations of the $\mathord{\sim}23$ Myr, V1298 Tau b revealed a H/He envelope fraction of $\mathord{\sim}40\%$ (Barat et al., 2024) similar to the predicted mass envelope fractions of TIC 434398831 b and c. While these simulations only account for mass-loss via photoevaporation, it is likely additional mass-loss mechanisms are at play and work in hand with photoevaporation to sculpt the distribution of the mature population (e.g. Gupta & Schlichting, 2019; Owen & Schlichting, 2023; Rogers et al., 2024).
@@ Acknowledgements
This work was enabled by observations made from the Gemini North telescope, located within the Maunakea Science Reserve and adjacent to the summit of Maunakea. We are grateful for the privilege of observing the Universe from a place that is unique in both its astronomical quality and its cultural significance. Observations in the paper made use of the High-Resolution Imaging instrument(s) Alopeke. Alopeke was funded by the NASA Exoplanet Exploration Program and built at the NASA Ames Research Center by Steve B. Howell, Nic Scott, Elliott P. Horch, and Emmett Quigley. Alopeke was mounted on the Gemini North telescope of the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation. on behalf of the Gemini partnership: the U.S. National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea).